Hurricanes strike Florida at an average rate of 1 every 5.3 years. What is the probability that Florida will be struck by two or more such hurricanes over the next decade
Answer #1

X ~ Poi ( )

Where \lambda = 1 per 5.3 years

So,

For 10 years,

= 10/5.3

Poisson probability distribution is

P(X) = e- * X / X!

So ,

P(X >= 2) = 1 - P(X <= 1)

= 1 - [ P(X = 0) + P(X = 1) ]

= 1 - [ e-(10/5.3) + e-(10/5.3) * (10/5.3) ]

= 0.5625